# What is Wrong with Perplexity for Long-context Language Modeling?

### v4

@@ 1 Introduction
Figure 1: (a) A constructed example to illustrate how LongPPL is calculated. We truncate the long context and calculate the generation probability difference (long-short difference, LSD, Eq. (2)) for each token based on the long and short contexts. A high LSD score indicates that the token’s generation is significantly enhanced by the long context, making it a key token in the long text. LongPPL is then obtained by calculating perplexity on these key tokens. (b) Long-context performance (LongBench (Bai et al., 2023b)) vs. perplexity measures (PPL and our LongPPL) computed on GovReport (Huang et al., 2021), a natural corpus. While PPL shows no correlation w.r.t. Longbench score, LongPPL achieves $-0.96$ Pearson correlation coefficient.
@@ 2.2 Extracting Key Tokens from Natural Texts
We evaluate the LSD score on LongEval, where we have knowledge of the key answer tokens. As shown in Figure 3(a), we compute the LSD score with a powerful long-context LLM, Mixtral-8x7B (Jiang et al., 2024), and find that answer tokens are clearly separated from the non-answer tokens: most answer tokens have LSD values higher than $2$, while most of the non-answer tokens concentrate around low LSD values (lower than $0.5$). When using LSD values alone to classify answer and non-answer tokens, we attain 85.6% accuracy (Figure 4(b)), indicating that LSD values are strongly indicative of the key tokens in long-context understandings.
From Figure 3(a), we find that a small proportion of non-answer tokens also have large LSDs (larger than $0.5$) and are thus confused together with key tokens. After analyzing, we find that these tokens can be further separated out by inspecting their Long-Context Likelihood (LCL) under long context:
@@
To summarize, we revisit our initial question why perplexity fails to represent long-context performance. As shown in Figure 4(a), most tokens in a natural corpus, GovReport (Huang et al., 2021), are long-context-irrelevant tokens with low LSD (lower than $0.5$), while only less than 10% tokens are highly influenced by long context (with LSD$>2$) and represent long-context abilities. Therefore, perplexity that averages over all tokens (Equation 1) does not represent the real long-context performance. Instead, combining the LSD (Equation 2) and the LCL (Equation 3) scores, we are able to accurately identify the answer tokens in LongEval with an accuracy of 98.2% (Figure 4(b)). Based on this result, in the next section, we design a new perplexity measure, LongPPL, that is tailored to reflect the long-context performance of LMs, by focusing on the key tokens.
@@ 3.2 Improving Long-context Capabilities with LongCE
Here, $\gamma>0$ is a hyper-parameter that sets a threshold on the maximal influence to avoid numerical instability. As a consequence of this reweighting term, too easy tokens (both short and long context give accurate prediction) and too hard tokens (neither short or long context predicts correctly) will have a weight around $1$, while those long-context-dependent tokens (high $P_{\theta}(x_{i}|{\bm{l}}_{i})$ and low $P_{\theta}(x_{i}|{\bm{s}}_{i})$) will be upweighted above $1$, proportionally to the context informativeness.
@@ 4.1 LongPPL Metric
PPL $I(x)\equiv 1$0.11 0.31 0.33
LongPPL-soft $I_{\rm soft}$ (Equation 7) -0.43 -0.21 -0.17
LongPPL-hard (default) $I$ (Equation 4) -0.96 -0.86 -0.84
@@
$\alpha=2,\beta=-2$ (default) -0.96 -0.86 -0.84
$\alpha=2,\beta=-1$0.96 -0.92 -0.92
$\alpha=1,\beta=-2$0.91 -0.73 -0.69
$\alpha=1,\beta=-1$0.97 -0.88 -0.87
@@ B.1 Detailed results of LongPPL
Mixtral-8x7B-32k 2.08 2.50 1.74 3.67
FILM-7B-32k 2.49 3.17 2.03 4.47
Mistral-7B-32k 2.68 3.49 2.19 4.25
Qwen1.5-14B-128k 2.97 2.93 2.33 5.23
Qwen2-7B-128k 2.99 2.73 2.29 4.97
Phi-3-small-128k 2.98 2.86 2.41 5.42
CLEX-7B-64k 3.70 4.60 2.92 4.13
Yi-6B-200k 3.62 3.92 2.86 5.11
Yarn-7B-128k 3.67 4.88 3.10 4.17
@@ B.3 Fine-grained results of LongCE
Task Domains Single-Document QA Multi-Document QA Summarization Few-shot Learning Code Completion Synthetic Tasks Avg.

### v5

@@ 1 Introduction
Figure 1: (a) A constructed example to illustrate how LongPPL is calculated. We truncate the long context and calculate the generation probability difference (long-short difference, LSD, Eq. (2)) for each token based on the long and short contexts. A high LSD score indicates that the token’s generation is significantly enhanced by the long context, making it a key token in the long text. LongPPL is then obtained by calculating perplexity on these key tokens. (b) Long-context performance (LongBench (Bai et al., 2023b)) vs. perplexity measures (PPL and our LongPPL) computed on GovReport (Huang et al., 2021), a natural corpus. While PPL shows no correlation w.r.t. Longbench score, LongPPL achieves 0.96-0.96- 0.96 Pearson correlation coefficient.
@@ 2.2 Extracting Key Tokens from Natural Texts
We evaluate the LSD score on LongEval, where we have knowledge of the key answer tokens. As shown in Figure 3(a), we compute the LSD score with a powerful long-context LLM, Mixtral-8x7B (Jiang et al., 2024), and find that answer tokens are clearly separated from the non-answer tokens: most answer tokens have LSD values higher than 222, while most of the non-answer tokens concentrate around low LSD values (lower than 0.50.50.5). When using LSD values alone to classify answer and non-answer tokens, we attain 85.6% accuracy (Figure 4(b)), indicating that LSD values are strongly indicative of the key tokens in long-context understandings.
From Figure 3(a), we find that a small proportion of non-answer tokens also have large LSDs (larger than 0.50.50.5) and are thus confused together with key tokens. After analyzing, we find that these tokens can be further separated out by inspecting their Long-Context Likelihood (LCL) under long context:
@@
To summarize, we revisit our initial question why perplexity fails to represent long-context performance. As shown in Figure 4(a), most tokens in a natural corpus, GovReport (Huang et al., 2021), are long-context-irrelevant tokens with low LSD (lower than 0.50.50.5), while only less than 10% tokens are highly influenced by long context (with LSD>2>2> 2) and represent long-context abilities. Therefore, perplexity that averages over all tokens (Equation 1) does not represent the real long-context performance. Instead, combining the LSD (Equation 2) and the LCL (Equation 3) scores, we are able to accurately identify the answer tokens in LongEval with an accuracy of 98.2% (Figure 4(b)). Based on this result, in the next section, we design a new perplexity measure, LongPPL, that is tailored to reflect the long-context performance of LMs, by focusing on the key tokens.
@@ 3.2 Improving Long-context Capabilities with LongCE
Here, $\gamma>0$ is a hyper-parameter that sets a threshold on the maximal influence to avoid numerical instability. As a consequence of this reweighting term, too easy tokens (both short and long context give accurate prediction) and too hard tokens (neither short or long context predicts correctly) will have a weight around 111, while those long-context-dependent tokens (high $P_{\theta}(x_{i}|{\bm{l}}_{i})$ and low $P_{\theta}(x_{i}|{\bm{s}}_{i})$) will be upweighted above 111, proportionally to the context informativeness.
@@ 4.1 LongPPL Metric
PPL $I(x)\equiv 1$0.18 0.24 0.27
LongPPL-soft $I_{\rm soft}$ (Equation 7) -0.43 -0.23 -0.19
LongPPL-hard (default) $I$ (Equation 4) -0.96 -0.90 -0.90
@@
$\alpha=2,\beta=-2$ (default) -0.96 -0.90 -0.90
$\alpha=2,\beta=-1$0.92 -0.94 -0.96
$\alpha=1,\beta=-2$0.94 -0.79 -0.75
$\alpha=1,\beta=-1$0.95 -0.92 -0.92
@@ B.1 Detailed results of LongPPL
Mixtral-8x7B-32k 1.99 2.33 1.70 3.59
FILM-7B-32k 2.28 2.81 1.95 4.35
Mistral-7B-32k 2.48 3.10 2.11 4.14
Qwen1.5-14B-128k 2.67 2.57 2.19 5.07
Qwen2-7B-128k 2.66 2.48 2.16 4.82
Phi-3-small-128k 2.66 2.58 2.28 5.29
CLEX-7B-64k 3.28 3.95 2.74 4.04
Yi-6B-200k 3.19 3.38 2.65 4.96
Yarn-7B-128k 3.47 4.51 2.98 4.06
@@ B.3 Fine-grained results of LongCE
Task Domains
Single-Document
QA
Multi-Document
QA
Summarization
Few-shot
Learning
Code
Completion
Synthetic
Tasks
Avg.